# Approximate SMT Counting Beyond Discrete Domains

### v1

@@ Abstract
We introduce $\mathsf{pact}$, an SMT model counter for hybrid formulas that uses hashing-based approximate model counting to estimate solutions with theoretical guarantees. $\mathsf{pact}$ makes a logarithmic number of SMT solver calls relative to the projection variables, leveraging optimized hash functions. $\mathsf{pact}$ achieves significant performance improvements over baselines on a large suite of benchmarks. In particular, out of 14,202 instances, $\mathsf{pact}$ successfully finished on 603 instances, while $\mathsf{Baseline}$ could only finish on 13 instances.
@@ I Introduction
In light of the availability of powerful SMT solvers, a natural next challenge is to explore the problem of model counting for SMT formulas. Despite demonstrated applications [4, 19], SMT counting remains underexplored. Most prior work has concentrated on problems where the underlying theory is discrete, such as bitvectors [20, 21, 22], linear integers  [23, 24, 25], and strings [26]. Recent studies have shown that in these cases, reducing the problem to Boolean model counting is often the most effective approach [27]. On the other hand, SMT solvers are widely employed for applications that require reasoning over both continuous and discrete variables. For example, to encode hybrid systems in cyber-physical systems or planning, it is essential to have both variables. The existing work in SMT counting is unable to handle SMT formulas with continuous variables.
In this work, we seek to remedy the aforementioned situation. In particular, we focus on a large class of SMT formulas, which we refer to as hybrid SMT formulas. The hybrid SMT formulas are defined over both discrete and continuous variables, and we are interested in solutions projected over discrete variables. Our investigations for the development of counting techniques for hybrid SMT formulas are motivated by their ability to model several interesting and relevant applications, such as robustness quantification of cyber-physical systems and counting reachable paths in software (for detailed discussion, see Section I-A).
The primary contribution of this work is the development of a model counting tool, $\mathsf{pact}$ [^1] A preliminary version of this paper appears at the Design Automation Conference (DAC) 2025. The source code is available at https://github.com/meelgroup/pact, for efficient projected counting of hybrid SMT formulas. The framework approximates the model count with $(\varepsilon,\delta)$ guarantees. $\mathsf{pact}$ supports SMT formulas with various theories, including linear and non-linear real numbers, floating-point arithmetic, arrays, bit-vectors, or any combination thereof. The projection variables are over bit-vectors. $\mathsf{pact}$ employs a hashing-based approximate model counting technique, utilizing various hash functions such as multiply-mod-prime, multiply-shift, and XOR. The algorithm makes $\mathcal{O}(\log(|S|))$ calls to the SMT oracle, where $S$ is the set of projection variables. We have implemented a user-friendly tool based on cvc5, which will be released post-publication. $\mathsf{pact}$ supports a diverse array of theories including QF_ABV, QF_BVFP, QF_UFBV, QF_ABVFPLRA, QF_ABVFP, QF_BVFPLRA.
To demonstrate runtime efficiency, we conduct an extensive empirical evaluation over 14,202 benchmark instances. Out of these 14,202 instances, $\mathsf{pact}$ successfully finished on 603 instances, while $\mathsf{Baseline}$ could finish only on 13 instances. Importantly, $\mathsf{Baseline}$ fails on counts above 3,570, while $\mathsf{pact}$ handles instances with over $1.7\times 10^{19}$ solutions.
@@ II-B Related Work
The success of propositional model counters, particularly approximate model counters, prompted efforts to extend the techniques to word-level constraints. Chistikov et al.[20] used bit-blasting to extend the propositional model counting technique to word-level benchmarks. Chakraborty et al.[21] designed $\mathsf{SMTApproxMC}$ by lifting the hash functions for word-level constraints. Kim and McCamant[22] designed a system to estimate model count of bitvector formulas. Ge et al. [30] developed a probabilistic polynomial-time model counting algorithm for bit-vector problems, and also developed a series of algorithms in the context of related SMT theories to compute or estimate [24, 31, 23] the number of solutions for linear integer arithmetic constraints. A closely related problem in the hybrid domain of Boolean and rational variables is Weighted Model Integration (WMI) [32], which involves computing the volume given the weight density over the entire domain. Extensive research addresses WMI through techniques such as predicate abstraction and All-SMT [33, 34], as well as methods leveraging knowledge compilation [35].
Hashing-based approximate model counting has been extensively studied over the past decades [36, 37, 38, 21, 39, 40, 20, 41, 42, 43, 44, 45]. Chakraborty et al.[1] showed that variations in a few key components in a generalized framework account for the diversity in prior approaches. While prior works focused on discrete domains such as Boolean variables[39, 40, 44] and bitvectors [21, 20], our approach extends this framework to hybrid SMT formulas.
@@ III Algorithm and Implementation
1:$\mathsf{L}\leftarrow\varnothing,\mathsf{it}\leftarrow 0$
2:$\mathsf{thresh},\mathsf{numIt},\ell\leftarrow\mathsf{GetConstants}(\varepsilon,\delta,\mathsf{family})$
3:$\mathsf{C}[0]\leftarrow\textsc{$\textstyle\mathsf{SaturatingCounter}$}(F,S,\mathsf{thresh})$
@@
1:$\mathsf{thresh}\leftarrow 1+9.84\left(1+\frac{\varepsilon}{1+\varepsilon}\right)\left(1+\frac{1}{\varepsilon}\right)^{2}$
@@
While each of the hashing constraints partitions the solution space into $p$ slices, we often want to partition it into more. To partition the solution space into $p^{c}$ cells, we use the Cartesian product of $c$ hash functions: $\mathcal{H}\times\mathcal{H}\times\dots\times\mathcal{H}$. We use the notation $H_{[i]}$ to denote the Cartesian product of the first $i+1$ hash functions, i.e., $H_{[i]}=h_{0}\times h_{1}\times\dots\times h_{i}$.
@@ III-F Implementation and Supported Theories
Logic Baseline $\mathsf{pact}_{\textit{prime}}$ $\mathsf{pact}_{\textit{shift}}$ $\mathsf{pact}_{\textit{xor}}$
QF_ABVFPLRA (30) -- -- -- 44 44
QF_ABVFP (333) 99 99 11 11 11 11 3131 3131
QF_ABV (2922) -- -- -- 287287 287287
QF_BVFPLRA (55) -- -- -- 3737 3737
QF_BVFP (10434) 22 22 7373 7373 $8$ $8$ 227227 227227
QF_UFBV (428) 22 22 99 99 22 22 1717 1717
Total (14202) 1313 1313 8383 8383 9191 9191 603603 603603
@@ IV Experimental Evaluation
We implemented the proposed algorithm on top of state-of-the-art SMT solver cvc5. We used the following experimental setup in the evaluation:
Baseline. As mentioned in Section I, the existing tools for SMT counting are unable to handle $\mathsf{Count{\mathcal{T}}_{\downarrow{\mathcal{P}}}}$ problems. The situation is similar to the early days of propositional model counting, wherein enumeration-based counters were employed as a baseline. In the same vein, we developed an enumeration-based counter, referred to as $\mathsf{Baseline}$, that uses the state-of-the-art SMT solver, cvc5, the same solver employed by $\mathsf{pact}$. $\mathsf{Baseline}$ operates by asking for a solution from the SMT solver and then appending a blocking clause that includes assignments to the projection variable. Subsequently, it asks the solver for another solution, continuing this process until the solver indicates UNSAT.
Benchmarks. Our benchmark suite comprises 14,202 instances from the SMT-Lib 2023 release. To minimize bias, we adopted a benchmark selection methodology inspired by early works on propositional model counting. We initially selected all instances supported by six theories. Subsequently, we filtered out instances where the number of solutions was very small (less than 500 models) or where even satisfiability was computationally challenging, as determined by cvc5’s inability to find a satisfying assignment within 5 seconds.
@@
    How does the runtime performance of $\mathsf{pact}$ compare to that of $\mathsf{Baseline}$, and how does the performance vary with different hash function families?
@@
Summary of Results. $\mathsf{pact}$ solves a significant number of instances from the benchmarks. It solved 603 instances, while the $\mathsf{Baseline}$ counted only 13 instances. Among different hash families, $\mathsf{pact}$ performs the best while it uses $\mathcal{H}_{\textit{xor}}$ hash functions. The accuracy of $\mathsf{pact}$ is also noteworthy; the average approximation error is 3.3% while using $\mathcal{H}_{\textit{xor}}$ hashes. [^2] The benchmarks and logfiles are available at https://doi.org/10.5281/zenodo.16413909
0100200300400500600700Benchmarks0500100015002000250030003500Runtime(s)$\mathsf{pact}_{\textit{xor}}$$\mathsf{pact}_{\textit{shift}}$$\mathsf{pact}_{\textit{prime}}$$\mathsf{Baseline}$
Figure 1: Comparison of time taken to count by $\mathsf{pact}$ and $\mathsf{Baseline}$.
@@ IV-A Performance of $\mathsf{pact}$
Instances solved. For each of the logic, we look at the number of instances solved. Out of 14,202 instances, $\mathsf{Baseline}$ could solve only 13 instances. Conversely, $\mathsf{pact}_{\textit{xor}}$ could solve 603 instances, demonstrating a substantial improvement compared to $\mathsf{Baseline}$. The performance varies across different logics, which we represent in Table I. The number of instances solved is $4.2\%$ of the total number of instances, which is expected, given the target problem for these instances is satisfiability, and $\mathsf{pact}$ solves a more complex problem.
Comparison of Hash Functions. In the Table I, we compare the performance of $\mathsf{pact}$ when it utilizes different hash functions for partitioning the solution space. While all of them perform much better in comparison to $\mathsf{Baseline}$, the best performance is shown by $\mathsf{pact}_{\textit{xor}}$, which solved 603 instances. The performances of $\mathsf{pact}_{\textit{prime}}$ and $\mathsf{pact}_{\textit{shift}}$ are similar, solving 83 and 91 instances.
[…]
Figure 2: Comparison of $\mathsf{Baseline}$ and $\mathsf{pact}$. Scalability w.r.t. the number of solutions of the formula.
Solving time comparison. A performance evaluation of $\mathsf{Baseline}$ and $\mathsf{pact}$ is depicted in Figure 1, which is a cactus plot comparing the solving time. The $x$-axis represents the number of instances, while the $y$-axis shows the time taken. A point $(i,j)$ in the plot represents that a solver solved $j$ benchmarks out of the 14,202 benchmarks in the test suite in less than or equal to $j$ seconds. The different curves plot the performance of $\mathsf{Baseline}$ and $\mathsf{pact}$ with different hash functions.
Performance against the number of solutions. The primary constraint in the enumeration-based method is its capacity for solution counting. The baseline model caps at 3,570 solutions, while $\mathsf{pact}$ counts up to $1.7\times 10^{19}$. Figure 2 presents a time versus solution count comparison, where the $x$-axis represents sorted instances by solution count and the $y$-axis indicates the time required for $\mathsf{Baseline}$ and $\mathsf{pact}$ to solve an instance. Thus, a point $(x,y_{1})$ for $\mathsf{Baseline}$ and $(x,y_{2})$ for $\mathsf{pact}$ represents the time taken to solve an instance with $x$ solutions, where $y_{1}$ seconds is the time for $\mathsf{Baseline}$ (exact count) and $y_{2}$ seconds is the time for $\mathsf{pact}$ (approximate count by $\mathsf{pact}$). A marker at $y=3600$ denotes a timeout. The graph demonstrates that $\mathsf{pact}$ solves many high-count instances rapidly, unlike $\mathsf{Baseline}$, which often times out as solution counts increase. The performance of $\mathsf{pact}_{\textit{xor}}$ is better than $\mathsf{pact}_{\textit{shift}}$ or $\mathsf{pact}_{\textit{prime}}$ in this regard as well - the maximum count returned by them are in magnitudes of $10^{7}$, while $\mathsf{pact}_{\textit{xor}}$ is of $10^{19}$.
@@ IV-B Quality of Approximation
From our benchmark set, only 13 instances were solved by $\mathsf{Baseline}$. To increase the number of instances for which we know the exact count, we also include the benchmarks with model counts between 100 and 500 in this section of the paper - resulting in 64 instances. We quantify the quality of approximation with the parameter error $e=\max\left(\frac{b}{s},\frac{s}{b}\right)-1$, where $b$ is the count from $\mathsf{Baseline}$ and $s$ from $\mathsf{pact}$. this definition of error aligns with the $\varepsilon$ used in the algorithm’s theoretical guarantees and can be interpreted as the observed value of $\varepsilon$. Analysis of all 64 cases found that for $\mathsf{pact}_{\textit{xor}}$, the maximum $e$ to be 0.260.26 0.260.26 and the average to be 0.030.03 0.030.03, signifying $\mathsf{pact}$ substantially outperforms its theoretical bounds, which is 0.8. In Figure 3, we illustrate the quality of approximation for these instances. The $x$-axis lists the instances, while the $y$-axis displays the relative error exhibited by a configuration of $\mathsf{pact}$. A dot $(x,y)$ in the graph indicates $x^{th}$ instance showed a relative error of $y$. The graph indicates that, for most instances, the error lies below 0.2, with a few instances falling between 0.2 and 0.8. The error for $\mathsf{pact}_{\textit{shift}}$ and $\mathsf{pact}_{\textit{prime}}$ is relatively higher than $\mathsf{pact}_{\textit{xor}}$, with average error being 0.070.07 0.070.07 and 0.120.12 0.120.12 and maximum error being 0.390.39 0.390.39 and 0.480.48 0.480.48. Our findings underline $\mathsf{pact}$’s accuracy and potential as a dependable tool for various applications.
@@ Acknowledgement
We are thankful to Jaroslav Bendik, Supratik Chakraborty, Ashwin Karthikeyan, Aina Niemetz, Mathias Preiner, Uddalok Sarkar, Mate Soos and Jiong Yang for the many useful discussions. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [RGPIN-2024-05956]. Part of the work was done when Arijit Shaw was a visiting graduate student at the University of Toronto. Computations were performed on the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto.

### v2

# arXiv:2507.18612v2[cs.LO] 26 Feb 2026

@@ Abstract
We introduce $\mathsf{pact}$, an SMT model counter for hybrid formulas that uses hashing-based approximate model counting to estimate solutions with theoretical guarantees. $\mathsf{pact}$ makes a logarithmic number of SMT solver calls relative to the projection variables, leveraging optimized hash functions. $\mathsf{pact}$ achieves significant performance improvements over baselines on a large suite of benchmarks. In particular, out of 3,119 instances, $\mathsf{pact}$ successfully finished on 456 instances, while the current state-of-the-art counter could only finish on 83 instances. [^1] A preliminary version of this paper appears at the Design Automation Conference (DAC) 2025.
@@ I Introduction
In light of the availability of powerful SMT solvers, a natural next challenge that has emerged is exploration of techniques for model counting for SMT formulas. A significant effort has been devoted to problems where the underlying theory is discrete, such as bitvectors [13, 31], linear integers  [25, 23, 27], and strings [3]. Recent studies have shown that in these cases, reducing the problem to Boolean model counting is often the most effective approach [41]. There are, however, several applications that require reasoning over both continuous and discrete variables, which we refer to as hybrid SMT formulas.
The hybrid SMT formulas are defined over both discrete and continuous variables, and we are interested in solutions projected over discrete variables. Our investigations for the development of counting techniques for hybrid SMT formulas are motivated by their ability to model several interesting and relevant applications, such as robustness quantification of cyber-physical systems and counting reachable paths in software (for detailed discussion, see Section I-A). The most relevant work to our setting is that of Chistikov, Dimitrova, and Majumdar [18], who introduced the notion of measured theories, which provides a unifying definition for quantitative problems across discrete and continuous SMT theories, as well as their combinations. Building on this framework, they developed hashing-based approximation algorithms for model counting over hybrid SMT formulas. Their techniques, however, face scalability hurdles, as discussed in Section IV and consequently, the development of scalability techniques for hybrid SMT formulas has remained a formidable challenge.
The primary contribution of this work is an affirmative answer to the aforementioned challenge via the development of a model counting tool, $\mathsf{pact}$, for efficient projected counting of hybrid SMT formulas. The $\mathsf{pact}$ approximates the model count with $(\varepsilon,\delta)$ guarantees. $\mathsf{pact}$ supports SMT formulas with various theories, including linear and non-linear real numbers, floating-point arithmetic, arrays, bit-vectors, or any combination thereof. The projection variables are over bit-vectors. $\mathsf{pact}$ employs a hashing-based approximate model counting technique, utilizing various hash functions such as multiply-mod-prime, multiply-shift, and XOR. The algorithm makes $\mathcal{O}(\log(|S|))$ calls to the SMT oracle, where $S$ is the set of projection variables. We have implemented a user-friendly open-source tool based on cvc5. $\mathsf{pact}$ supports a diverse array of theories including QF_ABV, QF_BVFP, QF_UFBV, QF_ABVFPLRA, QF_ABVFP, QF_BVFPLRA.
To demonstrate runtime efficiency, we conduct an extensive empirical evaluation over 3,119 benchmark instances. On these instances, $\mathsf{pact}$ successfully finished on 456 instances, while the current state-of-the-art [18] could finish only on 83 instances.
@@ I-A Applications
I-B Organization
The rest of the paper is organized as follows: We introduce the preliminaries and related work in section II. In section III, we present an overview of our framework, $\mathsf{pact}$. We describe our experimental methodology and results in section IV. Finally, we conclude in section V.
@@ II-B Related Work
The success of propositional model counters, particularly approximate model counters, prompted efforts to extend the techniques to word-level constraints. The most relevant work in this direction is due to Chistikov, Dimitrova, and Majumdar [18]. Their algorithm supports projected counting over hybrid domains by treating the continuous variables as existentially quantified. They reduce model counting to a sequence of SMT queries. To obtain an approximation with a desired precision, these SMT queries contain multiple copies of the original SMT formula, and the hashing constraints are applied to the duplicated free variables.
In the following years, Chakraborty et al.[13] designed $\mathsf{SMTApproxMC}$ for counting over theory of bit-vector by lifting the hash functions for word-level constraints. Kim and McCamant[31] designed a system to estimate model count of bitvector formulas. Ge et al. [24] developed a probabilistic polynomial-time model counting algorithm for bit-vector problems, and also developed a series of algorithms in the context of related SMT theories to compute or estimate [23, 26, 25] the number of solutions for linear integer arithmetic constraints.
A closely related problem in the hybrid domain of Boolean and rational variables is Weighted Model Integration (WMI) [8], which involves computing the volume given the weight density over the entire domain. Extensive research addresses WMI through techniques such as predicate abstraction and All-SMT [36, 37], as well as methods leveraging knowledge compilation [32].
Hashing-based approximate model counting has been extensively studied over the past decades [1, 2, 9, 13, 14, 15, 28, 42, 43, 46, 47]. Chakraborty et al.[16] showed that variations in a few key components in a generalized framework account for the diversity in prior approaches. While prior works focused on discrete domains such as Boolean variables[14, 15, 46] and bitvectors [13, 18], our approach extends this framework to hybrid SMT formulas.
@@ III Algorithm and Implementation
[…]
@@
[…]
@@ III-F Implementation and Supported Theories
Logic $\mathsf{CDM}$ $\mathsf{pact}_{\textit{prime}}$ $\mathsf{pact}_{\textit{shift}}$ $\mathsf{pact}_{\textit{xor}}$
QF_ABVFPLRA (19) - - - $1$
QF_ABVFP (13) - $1$ $1$ $7$
QF_ABV (2713) $11$ - - $284$
QF_BVFPLRA (46) - - - $30$
QF_BVFP (129) $71$ $23$ $37$ $117$
QF_UFBV (199) $1$ $9$ $2$ $17$
Total (3119) $83$ $33$ $40$ $456$
@@ IV Experimental Evaluation
We implemented [^2] The source code is available at https://github.com/meelgroup/pact. the proposed algorithm on top of state-of-the-art SMT solver cvc5. We used the following experimental setup in the evaluation:
Baseline. We compared the performance of $\mathsf{pact}$ with the current state-of-the-art tool by Chistikov, Dimitrova, and Majumdar [18]. We refer to the tool as $\mathsf{CDM}$, after the initials of its authors.
Benchmarks. Our benchmark suite comprises 3,119 instances from the SMT-Lib 2023 release. We adopted a benchmark selection methodology inspired by early works on propositional model counting. We initially selected all instances supported by six theories. Subsequently, we filtered out instances where the number of solutions was very small (less than 500 models) or where even satisfiability was computationally challenging, as determined by cvc5’s inability to find a satisfying assignment within 5 seconds. To avoid overweighting multiple near-identical benchmarks, we also sample to at most five benchmarks per cluster (benchmarks whose file names and differ only in index-level details).
@@
    How does the runtime performance of $\mathsf{pact}$ compare to that of $\mathsf{CDM}$, and how does the performance vary with different hash function families?
@@
Summary of Results. $\mathsf{pact}$ solves a significant number of instances from the benchmarks. It counted 456 instances, while the current state-of-the-art counted only 83 instances. Among different hash families, $\mathsf{pact}$ performs the best while it uses $\mathcal{H}_{\textit{xor}}$ hash functions. The accuracy of $\mathsf{pact}$ is also noteworthy; the average approximation error is 3.3% while using $\mathcal{H}_{\textit{xor}}$ hashes [^3] The benchmarks and logfiles are available at https://doi.org/10.5281/zenodo.16413909.
[…]
Figure 1: Cactus plot comparing performances of $\mathsf{pact}$ and $\mathsf{CDM}$.
@@ IV-A Performance of $\mathsf{pact}$
Instances solved. For each of the logic, we look at the number of instances solved. Out of 3,119 instances, $\mathsf{CDM}$ could solve only 83 instances. Conversely, $\mathsf{pact}_{\textit{xor}}$ could solve 456 instances, demonstrating a substantial improvement compared to $\mathsf{CDM}$. The performance varies across different logics, which we represent in Table I. The number of instances solved is $14.6\%$ of the total number of instances, which is expected, given the target problem for these instances is satisfiability, and $\mathsf{pact}$ solves a more complex problem.
Comparison of Hash Functions. In the Table I, we compare the performance of $\mathsf{pact}$ when it utilizes different hash functions for partitioning the solution space. The best performance is shown by $\mathsf{pact}_{\textit{xor}}$, which solved 456 instances. The performances of $\mathsf{pact}_{\textit{prime}}$ and $\mathsf{pact}_{\textit{shift}}$ are similar, solving 33 and 40 instances.
Solving time comparison. A performance evaluation of $\mathsf{CDM}$ and $\mathsf{pact}$ is depicted in Figure 1, which is a cactus plot comparing the solving time. The $x$-axis represents the number of instances, while the $y$-axis shows the time taken. A point $(i,j)$ in the plot represents that a solver solved $j$ benchmarks out of the 3,119 benchmarks in the test suite in less than or equal to $j$ seconds. The different curves plot the performance of $\mathsf{CDM}$ and $\mathsf{pact}$ with different hash functions.
@@ IV-B Quality of Approximation
As none of the existing tool returns the exact count, to get the exact number of solutions, we developed an enumeration-based counter, referred to as $\mathsf{enum}$. Here, we use the state-of-the-art SMT solver, cvc5, the same solver employed by $\mathsf{pact}$. From our benchmark set, only 9 instances were solved by $\mathsf{enum}$. To increase the number of instances for which we know the exact count, we also include the benchmarks with model counts between 100 and 500 in this section of the paper - resulting in 64 instances. We quantify the quality of approximation with the parameter error $e=\max\left(\frac{b}{s},\frac{s}{b}\right)-1$, where $b$ is the count from $\mathsf{enum}$ and $s$ from $\mathsf{pact}$. this definition of error aligns with the $\varepsilon$ used in the algorithm’s theoretical guarantees and can be interpreted as the observed value of $\varepsilon$. Analysis of all 64 cases found that for $\mathsf{pact}_{\textit{xor}}$, the maximum $e$ to be $0.26$ and the average to be $0.03$, signifying $\mathsf{pact}$ substantially outperforms its theoretical bounds, which is 0.8. In Figure 2, we illustrate the quality of approximation for these instances. The $x$-axis lists the instances, while the $y$-axis displays the relative error exhibited by a configuration of $\mathsf{pact}$. A dot $(x,y)$ in the graph indicates $x^{th}$ instance showed a relative error of $y$. The graph indicates that, for most instances, the error lies below 0.2, with a few instances falling between 0.2 and 0.8. The error for $\mathsf{pact}_{\textit{shift}}$ and $\mathsf{pact}_{\textit{prime}}$ is relatively higher than $\mathsf{pact}_{\textit{xor}}$, with average error being $0.07$ and $0.12$ and maximum error being $0.39$ and $0.48$. Our findings underline $\mathsf{pact}$’s accuracy and potential as a dependable tool for various applications.
@@ Acknowledgement
We are thankful to Jaroslav Bendik, Supratik Chakraborty, Dmitry Chistikov, Rayna Dimitrova, Ashwin Karthikeyan, Aina Niemetz, Mathias Preiner, Uddalok Sarkar, Mate Soos and Jiong Yang for the many useful discussions. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [RGPIN-2024-05956]. Part of the work was done when Arijit Shaw was a visiting graduate student at the University of Toronto. Computations were performed on the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto.